OpenStudy (anonymous):

how do you solve this for elimination: 4x+3y=14 -2x+y=8 please show steps

OpenStudy (anonymous):

muliply the -2x+y=8 by 2 you get -4x+2y=16

OpenStudy (anonymous):

now add these together 4x+-4x = 0x 3y+2y =5y and 14+16=30

OpenStudy (anonymous):

you get 5y=30 y=6

OpenStudy (anonymous):

-2x+6=8 x=-1

OpenStudy (anonymous):

x=-1 and y=6 is your answer

OpenStudy (anonymous):

now add these together 4x+-4x = 0x 3y+2y =5y and 14+16=30 how could i write this on my paper

OpenStudy (anonymous):

4x+3y=14 +(-4x+2y=16)

OpenStudy (anonymous):

solve this equation for y 4x+3y=14